### v1

@@ Abstract
By adapting Large Language Models (LLMs) to domain-specific tasks or enriching them with domain-specific knowledge, we can fully harness the capabilities of LLMs. Nonetheless, a gap persists in achieving simultaneous mutual enhancement between the server’s LLM and the downstream clients’ Small Language Models (SLMs). To address this, we propose FedCoLLM, a novel and parameter-efficient federated framework designed for co-tuning LLMs and SLMs. This approach is aimed at adaptively transferring server-side LLMs knowledge to clients’ SLMs while simultaneously enriching the LLMs with domain insights from the clients. To accomplish this, FedCoLLM utilizes lightweight adapters in conjunction with SLMs, facilitating knowledge exchange between server and clients in a manner that respects data privacy while also minimizing computational and communication overhead. Our evaluation of FedCoLLM, utilizing various public LLMs and SLMs across a range of NLP text generation tasks, reveals that the performance of clients’ SLMs experiences notable improvements with the assistance of the LLMs. Simultaneously, the LLMs enhanced via FedCoLLM achieves comparable performance to that obtained through direct fine-tuning on clients’ data.
Keywords:
LLMs Federated Learning Parameter-Efficient
@@ 1 Introduction
  • Mutual Knowledge Transfer Between LLMs and SLMs. Optimizing both LLMs and SLMs in different parties establishes a positive feedback loop that allows both LLMs and SLMs to evolve continuously. Initially, the LLMs on the server can disseminate general knowledge and capabilities to the clients’ SLMs, which are then trained on domain-specific data for downstream applications. Subsequently, these domain-specific SLMs can contribute their industry-specific knowledge back to the server’s LLMs. This transfer of specialized knowledge serves to enhance the LLMs’ understanding and capabilities, broadening its scope and depth. Nevertheless, mutually enhancing the server’s LLMs and clients’ SLMs is rarely exploited in literature.
@@ 2.1 Knowledge Distillation
Knowledge distillation is a technique that has gained significant attention in recent years, as it enables the transfer of knowledge from a larger teacher model to a smaller student model. One of the early works in this area was proposed by [8], which introduced the concept of knowledge distillation and demonstrated its effectiveness in improving the performance of compressed models. Since then, numerous studies have built upon this foundation and explored various distillation strategies [7]. For instance,  [4, 11] improved response-based knowledge distillation, which let the student model directly mimic the final prediction of the teacher model.  [1] proposed FitNets, which focus on matching intermediate representations between the teacher and student models. Another notable work is the relational knowledge distillation approach introduced by [14], which captures pairwise relationships between outputs to enhance distillation efficiency. Different from one-way knowledge distillation between the teacher and student networks, Deep Mutual Learning (DML) [22] allows two networks to learn from each other through their predicted probability distributions during the training process. These studies have demonstrated the potential of knowledge distillation in various tasks, such as image classification, object detection, and natural language processing.
@@ 2.2 Federated Learning for Large Language Models
Parameter-Efficient Fine-Tuning (PEFT) techniques [9] provide a straightforward remedy to address the challenges of communication overhead and fine-tuning expenses in Federated Learning (FL) for Large Language Models (LLMs) [16]. Numerous investigations have extended the application of PEFT methods within the FL framework tailored for LLMs. Notable contributions include FedPETuning [23], Federated Adapter Tuning [3], and Federated Prompt Tuning [24]. These research outcomes suggest that FL clients, particularly those with constrained storage capacities like mobile devices, can significantly profit from the adoption of PEFT methods. These approaches facilitate the sharing of LLMs across various tasks while necessitating the retention of only a minimal set of parameters per task, effectively decreasing storage demands. Through the utilization of PEFT methods, FL clients can adeptly tailor LLMs to meet their unique requirements, all the while minimizing communication overhead and reducing fine-tuning costs.
@@ 3.2 FedCoLLM
Toward this goal, we (1) adopt lightweight LoRA modules as the bridge to transfer the knowledge between clients and the server, (2) leverage mutual knowledge distillation to transfer knowledge between the LLM and the aggregated SLM, and (3) employ secure aggregation to protect the privacy of the knowledge transfer process.
Specifically, we assume that clients and the server share a SLM $g_{\phi}$ parameterized by $\phi$. Each client $k$ inserts a small low-rank adapter parameterized by $\theta_{k}$ into its local SLM. We denote a client’s local SLM with the added $\theta$ as $g_{\phi+\theta}$. Instead of training a global SLM $\phi$, clients collaboratively train a global LoRA module $\theta$. Then, Eq.(1) can be reformulated as follows:
@@
The server inserts a small low-rank adapter parameterized by $\omega$ into its LLM $f_{\psi}$. We denote the server’s LLM $f_{\psi}$ with the added $\omega$ as $f_{\psi+\omega}$. The server conducts the mutual knowledge transfer between the LLM $f_{\psi+\omega}$ and the global SLM $g_{\phi+\theta}$ through supervised fine-tuning and mutual knowledge distillation based on the auxiliary dataset $\mathcal{D}^{a}$.
We formulate the losses of supervised fine-tuning $f_{\psi+\omega}$ and $g_{\phi+\theta}$ (denoted as $\mathcal{L}_{\text{FT}}^{f}$ and $\mathcal{L}_{\text{FT}}^{g}$) as follows:
@@ 3.4 Privacy-Preserving Analysis
4:$\mathcal{D}^{a}$: the auxiliary distillation dataset;
5:$\mathcal{D}^{k}$: local datasets for each client $k$;
6:$\eta_{\omega}$: the learning rate for optimizing LLM $f_{\psi+\omega}$;
7:$\eta_{\theta}$: the learning rate for optimizing SLM $g_{\phi+\theta}$.
8:$f_{\psi+\omega}$, $g_{\phi+\theta}$.
@@ 4.1 Setup
Evaluation Metrics. We evaluate the model performance of fine-tuned LLMs and SLMs on the QA datasets using Accuracy as the primary metric. It’s worth noting that in our experiments, all methods undergo zero-shot evaluation, and we use the lm-evaluation-harness package [6]. Additionally, to assess the communication efficiency of our framework, we measure the communication cost by tracking the number of transmitted parameters.
@@ 4.2 Performance Evaluations
We conduct experiments with three settings. The first setting (denoted as S1) involves one server-side GPT-2-Large LLM and four client-side GPT-2-Small SLMs, the second setting (denoted as S2) involves one server-side OPT-6.7B LLM and four client-side OPT-1.3B SLMs, and the third setting (denoted as S3) involves one server-side LLaMa2-7B LLM and four client-side LLaMa2-1.3B SLMs. Tables 1 demonstrate the performance comparisons of our approach against other baselines. The top sub-table and the bottom sub-table compare the performance of FedCoLLM against baselines on the server’s LLM and clients’ SLMs, respectively.
The top sub-table of Table 1 shows that FedCoLLM significantly outperforms Zero-Shot on the server’s LLM in the three settings. It also shows that FedCoLLM achieves comparable performance of the Centralized scenario. For example, in the CQA dataset, FedCoLLM achieves a relative improvement of 41% over Zero-Shot on GPT-2-Large LLM, 47% on OPT-6.7B LLM, and 66% on LLaMa2-7B LLM. Furthermore, FedCoLLM nearly equals Centralized performance, reaching 98% on GPT-2-Large LLM, 99% on OPT-6.7B LLM, and 97% on LLaMa2-7B LLM.
The bottom sub-table of Table 1 shows that FedCoLLM performs better than the Zero-Shot, Standalone, and FedAvg due to the assistance of the server’s LLM. For example, in the CQA dataset, FedCoLLM achieves a relative improvement of 6% over Standalone on GPT-2-Small SLM, 8% on OPT-1.3B SLM, and 4% on LLaMa2-1.3B SLM. Furthermore, FedCoLLM achieves a relative improvement of 3% over FedAvg on GPT-2-Small SLM, 5% on OPT-1.3B SLM, and 2% on LLaMa2-1.3B SLM.
@@
Table 1: We evaluate FedCoLLM using three settings. The first setting (denoted as S1) involves one server-side GPT-2-Large LLM and four client-side GPT-2-Small SLMs, the second setting (denoted as S2) involves one server-side OPT-6.7B LLM and four client-side OPT-1.3B SLMs, and the third setting (denoted as S3) involves one server-side LLaMa2-7B LLM and four client-side LLaMa2-1.3B SLMs. The top and bottom sub-tables compare the performance of FedCoLLM against baselines on the server’s LLM and clients’ SLMs, respectively. The results reported in the bottom sub-table are the average of all clients.
@@ 4.3 Communication Cost
Model Method Transmitted Param Size (MB) Param Percent (%)

### v2

@@ Abstract
By adapting Large Language Models (LLMs) to domain-specific tasks or enriching them with domain-specific knowledge, we can fully harness the capabilities of LLMs. Nonetheless, a gap persists in achieving simultaneous mutual enhancement between the server’s LLM and the downstream clients’ Small Language Models (SLMs). To address this, we propose FedCoLLM, a novel and parameter-efficient federated framework designed for co-tuning LLMs and SLMs. This approach is aimed at adaptively transferring server-side LLMs knowledge to clients’ SLMs while simultaneously enriching the LLMs with domain insights from the clients. To accomplish this, FedCoLLM utilizes lightweight adapters in conjunction with SLMs, facilitating knowledge exchange between server and clients in a manner that respects data privacy while also minimizing computational and communication overhead. Our evaluation of FedCoLLM, utilizing various public LLMs and SLMs across a range of NLP text generation tasks, reveals that the performance of clients’ SLMs experiences notable improvements with the assistance of the LLMs. Simultaneously, the LLMs enhanced via FedCoLLM achieves comparable performance to that obtained through direct fine-tuning on clients’ data. Our code has been contributed to the FATE open-source project and is now publicly accessible at https://github.com/FederatedAI/FATE-LLM/tree/main/python/fate_llm/algo/fedcollm.
@@ 1 Introduction
  • Mutual Knowledge Transfer Between LLMs and SLMs. Optimizing LLMs and SLMs separately forms a positive feedback loop for continuous evolution. Server-side LLMs first transfer general knowledge to client-side SLMs, which are then fine-tuned on domain-specific data for downstream tasks. These domain-specialized SLMs can in turn feed industry-specific knowledge back to server LLMs, improving their understanding and expanding their scope and depth. However, such mutual enhancement between server LLMs and client SLMs remains largely unexplored in existing literature.
@@ 2.1 Knowledge Distillation
Knowledge distillation is a technique that has gained significant attention in recent years, as it enables the transfer of knowledge from a larger teacher model to a smaller student model. One of the early works in this area was proposed by [9], which introduced the concept of knowledge distillation and demonstrated its effectiveness in improving the performance of compressed models. Since then, numerous studies have built upon this foundation and explored various distillation strategies [8]. For instance, two notable work  [4, 13] improved response-based knowledge distillation, which let the student model directly mimic the final prediction of the teacher model. FitNets [1] focus on matching intermediate representations between the teacher and student models. Another notable work is the relational knowledge distillation approach introduced by [16], which captures pairwise relationships between outputs to enhance distillation efficiency. Different from one-way knowledge distillation between the teacher and student networks, Deep Mutual Learning (DML) [23] allows two networks to learn from each other through their predicted probability distributions during the training process. These studies have demonstrated the potential of knowledge distillation in various tasks, such as image classification, object detection, and natural language processing.
@@ 2.2 Federated Learning for Large Language Models
Parameter-Efficient Fine-Tuning (PEFT) techniques [10] provide a straightforward remedy to address the challenges of communication overhead and fine-tuning expenses in Federated Learning (FL) for Large Language Models (LLMs) [7, 11, 6]. Numerous investigations have extended the application of PEFT methods within the FL framework tailored for LLMs. Notable contributions include FedPETuning [24], Federated Adapter Tuning [3], and Federated Prompt Tuning [25]. These research outcomes suggest that FL clients, particularly those with constrained storage capacities like mobile devices, can significantly profit from the adoption of PEFT methods. These approaches facilitate the sharing of LLMs across various tasks while necessitating the retention of only a minimal set of parameters per task, effectively decreasing storage demands. Through the utilization of PEFT methods, FL clients can adeptly tailor LLMs to meet their unique requirements, all the while minimizing communication overhead and reducing fine-tuning costs.
@@ 3.2 FedCoLLM
Toward this goal, we (1) adopt lightweight LoRA modules as the bridge to transfer the knowledge between clients and the server, (2) leverage mutual knowledge distillation to transfer knowledge between the LLM and the aggregated SLM, and (3) employ secure aggregation to protect the privacy of the knowledge transfer process. Specifically, we assume that clients and the server share a SLM $g_{\phi}$ parameterized by $\phi$. Each client $k$ inserts a small low-rank adapter parameterized by $\theta_{k}$ into its local SLM. We denote a client’s local SLM with the added $\theta$ as $g_{\phi+\theta}$. Instead of training a global SLM $\phi$, clients collaboratively train a global LoRA module $\theta$. Then, Eq.(1) can be reformulated as follows:
@@
The server inserts a small low-rank adapter parameterized by $\omega$ into its LLM $f_{\psi}$. We denote the server’s LLM $f_{\psi}$ with the added $\omega$ as $f_{\psi+\omega}$. The server conducts the mutual knowledge transfer between the LLM $f_{\psi+\omega}$ and the global SLM $g_{\phi+\theta}$ through supervised fine-tuning and mutual knowledge distillation based on the auxiliary dataset $\mathcal{D}^{a}$. We formulate the losses of supervised fine-tuning $f_{\psi+\omega}$ and $g_{\phi+\theta}$ (denoted as $\mathcal{L}_{\text{FT}}^{f}$ and $\mathcal{L}_{\text{FT}}^{g}$) as follows:
@@ 3.4 Privacy-Preserving Analysis
[…]
@@ 4.1 Setup
Evaluation Metrics. We evaluate the model performance of fine-tuned LLMs and SLMs on the QA datasets using Accuracy as the primary metric. It’s worth noting that in our experiments, all methods undergo zero-shot evaluation, and we use the lm-evaluation-harness package. Additionally, to assess the communication efficiency of our framework, we measure the communication cost by tracking the number of transmitted parameters.
@@ 4.2 Performance Evaluations
We conduct experiments with three settings. The first setting (denoted as S1) involves one server-side GPT-2-Large LLM and four client-side GPT-2-Small SLMs, the second setting (denoted as S2) involves one server-side OPT-6.7B LLM and four client-side OPT-1.3B SLMs, and the third setting (denoted as S3) involves one server-side LLaMa2-7B LLM and four client-side LLaMa2-1.3B SLMs. Tables 1 2 demonstrate the performance comparisons of our approach against other baselines. The top sub-table and the bottom sub-table compare the performance of FedCoLLM against baselines on the server’s LLM and clients’ SLMs, respectively.
The Table 1 shows that FedCoLLM significantly outperforms Zero-Shot on the server’s LLM in the three settings. It also shows that FedCoLLM achieves comparable performance of the Centralized scenario. For example, in the CQA dataset, FedCoLLM achieves a relative improvement of 41% over Zero-Shot on GPT-2-Large LLM, 47% on OPT-6.7B LLM, and 66% on LLaMa2-7B LLM. Furthermore, FedCoLLM nearly equals Centralized performance, reaching 98% on GPT-2-Large LLM, 99% on OPT-6.7B LLM, and 97% on LLaMa2-7B LLM.
The Table 2 shows that the SLM of FedCoLLM performs better than the Zero-Shot, Standalone, and FedAvg due to the assistance of the server’s LLM. For example, in the CQA dataset, FedCoLLM achieves a relative improvement of 6% over Standalone on GPT-2-Small SLM, 8% on OPT-1.3B SLM, and 4% on LLaMa2-1.3B SLM. Furthermore, FedCoLLM achieves a relative improvement of 3% over FedAvg on GPT-2-Small SLM, 5% on OPT-1.3B SLM, and 2% on LLaMa2-1.3B SLM.
@@
Table 1: We evaluate the server-side LLM performance of FedCoLLM using three experimental settings. Setting 1 (denoted as S1) deploys one server-side GPT-2-Large LLM paired with four client-side GPT-2-Small SLMs; Setting 2 (denoted as S2) uses one server-side OPT-6.7B LLM and four client-side OPT-1.3B SLMs; and Setting 3 (denoted as S3) adopts one server-side LLaMa2-7B LLM with four client-side LLaMa2-1.3B SLMs.
@@
Table 2: We evaluate the client-side SLM performance of FedCoLLM across three experimental settings. The results reported in the table are the average of all clients.
@@ 4.3 Communication Cost
Model Method Transmitted Param Size (M) Param Percent (%)